OpenStudy (anonymous):

find formula for the inverse of function f(x)=sqrt{10-3x}

OpenStudy (amorfide):

let y=f(x) y=sqrt(10-3x) you want to make x the subject square both sides y²=10-3x 3x=10-y² x=(10-y²)/3 now replace y with x f^-1(x)=(10-x²)/3

OpenStudy (anonymous):

\[f(x)=y=\sqrt{10-3x}\]To find the inverse first we need to exchange the variable's place.\[x=\sqrt{10-3y}\]Now we need to solve for y\[x^{2}=10-3y ...squaring... both ..side\]\[3y=10-x^{2}\]\[f^{-1}(x)=y=\frac{ 10-x^{2} }{ 3 }\]

OpenStudy (anonymous):

the Answere is f-1(x)=-1/3x2+10/3

OpenStudy (anonymous):

You have a big problem in algebra. What is the d/c b/n we wright above and what u just posted?

OpenStudy (amorfide):

our answers are the same as your answer mahdi

OpenStudy (anonymous):

\[f^{-1} (x)=-\frac{ 1 }{ 3} x^{2}+\frac{ 10 }{ 3 }\]

OpenStudy (amorfide):

that is the same as our answer except you made it 2 fractions, we left it as one